# MIBench: A Comprehensive Benchmark for Model Inversion Attack and Defense

### v2

@@ A.4 Details of Attack Process
This step generate random latent vectors and distributes them for each attack target. Most attack methods use a random distribution strategy. Mirror, PPA and BREP calculate the score of each latent vector corresponding to each label, and for each label the vectors with the highest scores are selected.
@@
The previous step distributes latent vectors for each label, and this step optimizes the latent vectors by calculating the scores of the latent vectors corresponding to the labels and selecting the few vectors with the highest scores. Although currently there are no algorithms use this step, it can be added into the attack algorithms that use conditional generators, e.g. PLGMI (Yuan et al., 2023a) and LOKT (Nguyen et al., 2023a).
@@
Optimization is the key step for the attack process, it accepts the initialized latent vectors and attack labels as input and outputs the optimized inverted images. We provide several kinds of optimization strategies of each attack method as follows.
@@
Variance White-Box Optimization. It optimizes results from a Gaussian distribution of latent vectors corresponding to the target labels, and images are generated by random sampling from the optimized distribution. It is used by KEDMI (Chen et al., 2021b).
@@
BREP Optimization. A specific optimizer for BREPMI algorithm (Kahla et al., 2022). It uses a boundary repelling strategy for gradient simulation.
@@
Reinforcement Learning Optimization. Optimizing the latent vectors via reinforcement learning. Used by RLB attack method (Han et al., 2023).
@@
This step works by calculating the scores of each image and selecting the part of the image with the highest score as the result of the attack. It is used by PPA (Struppek et al., 2022).
@@
A.5 Details of Evaluation
We provide the following four evaluation metrics to evaluate the effectiveness of the attack.
1.
Classification Accuracy. The metric uses a given classifier to classify the inverted images and measures the top- and top- accuracy for target labels. The higher the reconstructed samples achieve attack accuracy on another classifier trained with the same dataset, the more private information in the dataset can be considered to be exposed (Zhang et al., 2020a).
2.
Feature Distance. The feature is defined as the output of the given classifier’s penultimate layer. We compute the shortest feature distance between inverted images and private samples for each class and calculate the average distance. Smaller feature distance means more similar features to the private image.
- 3.
- 4.
Appendix B Experimental Details
@@
B.1 Experiment Settings
We conducted experiments at both low and high resolution scenarios. For the low resolution experiments, we employed classifiers with a resolution of as the target models, and an ResNet- with a resolution of served as the evaluation model. In the high resolution experiments, we used classifiers with a resolution of as the target models, with an Inception-v3 model having a resolution of as the evaluation model. Additionally, the computation resources utilized in our experiments including NVIDIA RTX 4090 and NVIDIA A100.
@@ B.3 Classifiers
| CelebA | VGG-16 | 0.8826 | |
| ResNet-152 | 0.9371 | ||
| ResNet-50 | 0.9588 | ||
| ResNet-152 | 0.9003 | ||
| Inception-v3 | 0.9216 | ||
| FaceScrub | VGG-16 | 0.8785 | |
| ResNet-152 | 0.9825 | ||
| ResNet-50 | 0.9938 | ||
| ResNet-152 | 0.9225 | ||
| ResNeSt-101 | 0.9329 | ||
| Inception-v3 | 0.9445 |
@@
| TL |
B.4 Evaluation.
The definitions of the evaluation metrics are detailed in Section A.5. Here, we present the specific details of the metrics used in the experiments.
@@
Appendix C More Experimental Results
@@ C.1 Sample Diversity
Following the settings of Section 4.2 and 4.3, we computed the Precision-Recall (Kynkäänniemi et al., 2019) and Density-Coverage (Kynkäänniemi et al., 2019) to evaluate the intra-class diversity for each attack method. The results are presented in Table 9, 10, 11 and 12. It tends to be that the attacks with stronger GAN priors get higher scores, such as Mirror, C2FMI and PPA.
@@ C.2 Evaluation on Different Target Classifiers
In addition to attacking IR-152 and ResNet-152 in Section 4.2 and 4.3, we extend our experiments on more different target classifiers. We evaluate attacks on VGG- (Simonyan & Zisserman, 2014) and ResNeSt- (Zhang et al., 2022a) in low and high resolution settings, respectively. The results are presented in Table 13 and 14.
@@
Besides the aforementioned CNN-based classifiers, we further analyze MI attacks on ViT-based models. Table 15 presents further experiments on the MaxViT (Tu et al., 2022) under the high resolution (224×224) scenario. Other experimental settings are continuous with the main paper, with FFHQ as the public dataset and FaceScrub as the private dataset. The test accuracy of the target MaxViT is 94.61%.
@@ C.3 Evaluation on More Combination of Datasets
Evaluations in the Section 4 are conducted under the same dataset combination of FFHQ as the public dataset and FaceScrub as the private dataset. Therefore, we design more combination of datasets in this part to further assess the transferability of different attacks. The results are listed in Table 16 and 17. The visual results are shown in Figure 6 and 7.
Except for the typical face classification task, we have conducted more experiments on the dog breed classification task under the high resolution (224×224) scenario, which includes two non-facial datasets, Stanford Dogs (Dataset, 2011) and Animal Faces-HQ Dog (AFHQ) (Choi et al., 2020). The public dataset is AFHQ while the private dataset is Stanford Dogs. The target model is ResNet-152 of 77.45% test accuracy, which follows the same setting in PPA. Table 18 lists the evaluation results.
@@
| IF-GMI |
C.4 Evaluation on LS Defense Method
To evaluate the LS defense method, we follow the settings in the official paper (Struppek et al., 2023) and evaluate it in more settings, shown in Table 19. To better evaluate the effectiveness of LS defense algorithms by making the undefended classifiers slightly less accurate than the defense-imposed classifiers through an early-stop strategy when training.
| Setting | Defense | Test Acc | ||||
| A | - | |||||
| LS | ||||||
| B | - | |||||
| LS |
C.5 Validation for Defense Methods on Models with Low Accuracy
| Method | Hyperparameters | Test Acc | |||||
| NO Defense | - | 92.170 | 0.686 | 0.914 | 262.471 | 0.767 | 68.454 |
| MID | 88.240 | 0.568 | 0.820 | 246.021 | 0.757 | 69.663 | |
| BiDO | 88.620 | 0.582 | 0.874 | 275.453 | 0.793 | 68.248 | |
| TL | 89.160 | 0.316 | 0.616 | 279.439 | 0.897 | 63.241 |
C.6 Evaluation on Different Loss Functions
In recent years, various attack algorithms have attempted to mitigate the effects of gradient vanishing by employing different loss functions. In this part, we investigate the impact of identity loss functions on the success rate of model inversion attacks. Specifically, we adopt PPA with FFHQ prior to attack a ResNet-152 classifier pre-trained on FaceScrub. The comparison of the results is presented in Table 21. Our findings indicate that the Poincare loss function yields the highest performance without model augmentation, whereas the Logit loss function achieves the best results with model augmentation.
@@ C.6 Evaluation on Different Loss Functions
C.7 More evaluation on VMI, RLBMI and PPA
Considering the high computational overhead of RLBMI and VMI, we only experiment at a low resolution settings for classes. The public and private datasets are different part of CelebA dataset. The results are shown in Table 22.
@@ C.7 More evaluation on VMI, RLBMI and PPA
Appendix D Limitations and Future Plans
Our benchmark mainly focuses on GAN-based MI attacks and MI defenses applied in the classifier training stage. We will extend our benchmark to wider range of MI methods, including learning-based MI attacks and MI defenses applied to the classifier output. Furthermore, the concept of MI also pervades across modalities (e.g. text (Parikh et al., 2022; Zhang et al., 2023; Carlini et al., 2019; 2021; Yu et al., 2023; Nasr et al., 2023) and graph learning (Zhang et al., 2021; 2022b; Zhou et al., 2023; Wu et al., 2022; He et al., 2021)) beyond computer vision domain, where our benchmark concentrates in current stage. Expansion to new modalities is a promising direction for our benchmark to further explore the privacy threats in other fields, leading to more generalization in the AI security. In addition to developing new algorithms, it is also essential to conduct further research on the MI attacks and defenses to make in-depth analysis about their characteristics and bring valuable new insights.

### v3

@@ A.4 Details of Attack Process
This step generates random latent vectors and distributes them for each attack target. Most attack methods use a random distribution strategy. Mirror, PPA, and BREP calculate the score of each latent vector corresponding to each label, and for each label,l the vectors with the highest scores are selected.
@@
The previous step distributes latent vectors for each label, and this step optimizes the latent vectors by calculating the scores of the latent vectors corresponding to the labels and selecting the few vectors with the highest scores. Although currently there are no algorithms that use this step, it can be added into the attack algorithms that use conditional generators, e.g., PLGMI [60] and LOKT [34].
@@
Optimization is the key step for the attack process. It accepts the initialized latent vectors and attack labels as input and outputs the optimized inverted images. We provide several kinds of optimization strategies of each attack method as follows.
@@
Variance White-Box Optimization. It optimizes results from a Gaussian distribution of latent vectors corresponding to the target labels, and images are generated by random sampling from the optimized distribution. It is used by KEDMI [3].
@@
BREP Optimization. A specific optimizer for BREPMI algorithm [21]. It uses a boundary repelling strategy for gradient simulation.
@@
Reinforcement Learning Optimization. Optimizing the latent vectors via reinforcement learning. Used by RLB attack method [16].
@@
This step works by calculating the scores of each image and selecting the part of the image with the highest score as the result of the attack. It is used by PPA [43].
@@
Appendix B Evaluation Details
@@
B.1 Evaluation Settings
Following the standard settings in existing works, we conducted experiments at both low and high resolution scenarios. For the low-resolution experiments, we employed classifiers with a resolution of as the target models, and a ResNet- with a resolution of served as the evaluation model. In the high resolution experiments, we used classifiers with a resolution of as the target models, with an Inception-v3 model having a resolution of as the evaluation model. Additionally, the computation resources utilized in our experiments including NVIDIA RTX 4090 and NVIDIA A100.
@@ B.3 Classifiers
| CelebA | VGG-16 | 0.882 | |
| ResNet-152 | 0.937 | ||
| ResNet-50 | 0.958 | ||
| ResNet-152 | 0.900 | ||
| Inception-v3 | 0.921 | ||
| FaceScrub | VGG-16 | 0.878 | |
| ResNet-152 | 0.982 | ||
| ResNet-50 | 0.993 | ||
| ResNet-152 | 0.922 | ||
| ResNeSt-101 | 0.932 | ||
| Inception-v3 | 0.944 |
@@
| TL | ||
| TTS | ||
| Purifier | ||
| SSD |
B.4 Evaluation Metrics
The definitions of the evaluation metrics are detailed in Section 3.3. Here, we present the specific details of the metrics used in the experiments.
@@ B.4 Evaluation Metrics
Appendix C More Evaluation
@@ C.1 Sample Diversity
Following the settings of Section 4.2 and 4.3, we computed the Precision-Recall [27] and Density-Coverage [27] to evaluate the intra-class diversity for each attack method. The results are presented in Table 10, 11, 12 and 13. It tends to be that the attacks with stronger GAN priors get higher scores, such as Mirror, C2FMI and PPA.
@@ C.2 Evaluation on Different Target Classifiers
In addition to attacking IR-152 and ResNet-152 in Section 4.2 and 4.3, we extend our experiments on more different target classifiers. We evaluate attacks on VGG- [41] and ResNeSt- [62] in low- and high-resolution settings, respectively. The results are presented in Table 14 and 15.
@@
Besides the aforementioned CNN-based classifiers, we further analyze MI attacks on ViT-based models. Table 16 presents further experiments on the MaxViT [48] under the high resolution (224×224) scenario. Other experimental settings are continuous with the main paper, with FFHQ as the public dataset and FaceScrub as the private dataset. The test accuracy of the target MaxViT is 94.61%.
@@ C.3 Evaluation on More Combination of Datasets
Evaluations in the Section 4 are conducted under the same dataset combination of FFHQ as the public dataset and FaceScrub as the private dataset. Therefore, we design more combination of datasets in this part to further assess the transferability of different attacks. The results are listed in Table 17 and 18. The visual results are shown in Figure 9 and 10.
Except for the typical face classification task, we have conducted more experiments on the dog breed classification task under the high resolution (224×224) scenario, which includes two non-facial datasets, Stanford Dogs [5] and Animal Faces-HQ Dog (AFHQ) [4]. The public dataset is AFHQ while the private dataset is Stanford Dogs. The target model is ResNet-152 of 77.45% test accuracy, which follows the same setting in PPA. Table 19 lists the evaluation results.
@@
| IF-GMI | 48.972 |
C.4 Evaluation on Different Loss Functions
In recent years, various attack algorithms have attempted to mitigate the effects of gradient vanishing by employing different loss functions. In this part, we investigate the impact of identity loss functions on the success rate of model inversion attacks. Specifically, we adopt PPA with FFHQ prior to attack a ResNet-152 classifier pre-trained on FaceScrub. The comparison is presented in Table 20. Our findings indicate that the Poincaré loss function yields the highest performance without model augmentation, whereas the Logit loss function achieves the best results with model augmentation, serving as the optimal loss choices for future studies.
@@ C.4 Evaluation on Different Loss Functions
C.5 Evaluation on Time Overhead
We conduct an extra evaluation on the time overhead among all the MI attacks. The results are recorded in the Table 21. The time complexities of most algorithms fall within the same order of magnitude, which is around 100 seconds per class. Notably, the time cost of VMI and RLBMI is extraordinarily higher than other attacks, indicating that both methods are hard to evaluate under the identical scenarios in the main evaluation and relatively unpractical in real-world settings. Therefore, we put the relevant experiments on the next section (Supp.C.6).
| Attacks | |
| GMI | 78 |
| KEDMI | 73 |
| VMI | 5910 |
| Mirror(white-box) | 43 |
| PPA | 124 |
| PLGMI | 59 |
| LOMMA+GMI | 133 |
| LOMMA+KEDMA | 127 |
| IF-GMI | 83 |
| Mirror(black-box) | 161 |
| C2FMI | 83 |
| RLBMI | 10633 |
| BREPMI | 298 |
| LOKT | 219 |
C.6 More evaluation on VMI, RLBMI and PPA
Considering the high computational overhead of RLBMI and VMI, we only experiment at a low resolution settings for classes. The public and private datasets are splitted part of CelebA dataset. The results are shown in Table 22.
@@ C.6 More evaluation on VMI, RLBMI and PPA
C.7 More Visualization
We visualize more results in Figure 6, 9 and 10. The target model is ResNet- trained on CelebA dataset in high resolution setting.
Appendix D Special Experiments
D.1 Validation on LS Defense Method
To validate the correctness of our implemented LS defense, we follow the settings in the official paper [44] and evaluate it in more settings, shown in Table 24. To better evaluate the effectiveness of LS defense algorithms, we make the undefended classifiers slightly less accurate than the defense-imposed classifiers through an early-stop strategy when training.
| Setting | Defense | Test Acc | ||||
| A | - | |||||
| LS | ||||||
| B | - | |||||
| LS |
D.2 Validation for Defense Methods on Models with Low Accuracy
| Method | Hyperparameters | Test Acc | |||||
| NO Defense | - | 92.170 | 0.686 | 0.914 | 262.471 | 0.767 | 68.454 |
| MID | 88.240 | 0.568 | 0.820 | 246.021 | 0.757 | 69.663 | |
| BiDO | 88.620 | 0.582 | 0.874 | 275.453 | 0.793 | 68.248 | |
| TL | 89.160 | 0.316 | 0.616 | 279.439 | 0.897 | 63.241 |
D.3 The abnormal visualization of Mirror
To determine the cause of visualization anomalies in the Mirror method, we conduct a simple ablation experiment on its optimization objective, as denoted in Table 26 and Fig 8. The experimental results demonstrate that when substituting the optimization objective with the space latent vectors, the visualization of Mirror returns to normality. Therefore, the underlying reason of the visualization anomalies is likely attributed to genetic-algorithm-induced corruption of StyleGAN’s latent vectors. As a disentangled latent vector that deeply participates in StyleGAN’s hierarchical progressive image generation process, the compromised latent vector ultimately leads to image artifacts in the generated outputs.
| Scenario | Space | ||||
| White-box | |||||
| Black-box | |||||
D.4 User Study
In this section, we conduct a user study considering a subjective evaluation of MI attacks to further evaluate their effectiveness. We use different attack results in the resolution scenarios where the public dataset is FFHQ and the private dataset is FaceScrub. We randomly select classes for user study and publish survey questionnaires on the credamo website. The user interface is shown in Fig. 7. For each class, users are shown private images (above in Fig.7) as examples and select the reconstructed images (below in Fig.7) that have the top-3 visual similarity to the examples. In total, users have participated in the annotation, and the result is shown in Table 27.
| Attacks | |
| GMI | 0.035 |
| KEDMI | 0.084 |
| Mirror(white-box) | 0.493 |
| PPA | 0.564 |
| PLGMI | 0.326 |
| LOMMA+GMI | 0.074 |
| LOMMA+KEDMA | 0.083 |
| IF-GMI | 0.419 |
| BREPMI | 0.026 |
| Mirror(black-box) | 0.027 |
| C2FMI | 0.066 |
| LOKT | 0.059 |
Appendix E Limitations and Future Plans.
During the writing process, there were new works about MI published in the top-tier conferences, such as [38]. As a toolbox for MI research, our benchmark is continuously updated and all the new works are scheduled to be implemented in the near future. In addition to developing new algorithms, it is also essential to conduct further research on the MI attacks and defenses to make in-depth analysis about their characteristics and bring valuable new insights.